Assume that your filing status is single, you are not head of household, and you are...

  1. Assume that your filing status is single, you are not head of household, and you are 35 years old. Compute your refund based on the following data: Your wages are $65,000, and you had $15,000 withheld for federal taxes and $3,000 for state taxes. What is your refund? Answer:__________________
  2. Compute your refund using the same data as in a. except you are married and filing jointly, and your spouse is also 35 years old. Answer:_____________
  3. Assume the same facts as in b. except you own your own home on which you paid $15,000 in interest and $5,000 in real estate taxes during the year. What is your refund? Answer:____________________
  4. Using the same facts as in c. assume that you and your spouse both contributed $2,000 to a traditional IRA. Compute your refund. Answer:____________
Answer #1

2017 Federal Tax rates are as below:

Single
Taxable Income Tax Rate
$0 – $9,325 10%
$9,326 – $37,950 $932.50 plus 15% of the amount over $9,325
$37,951 – $91,900 $5,226.25 plus 25% of the amount over $37,950
$91,901 – $191,650 $18,713.75 plus 28% of the amount over $91,900
$191,651 – $416,700 $46,643.75 plus 33% of the amount over $191,650
$416,701 – $418,400 $120,910.25 plus 35% of the amount over $416,700
$418,401 or more $121,505.25 plus 39.6% of the amount over $418,400

As per question wages is $65,000, so it would fall in the 3rd slab marked in bold above.

Tax will be $5,226.25 plus 25% of the amount over $37,950.

= 5,226.25 + (65000-37951)*0.25 = 5,226.25+ 6762.25= $11,988.50.

Withheld for Federal Taxes = $15,000

Refund = $15,000 - $11,988.50= $3011.5

Rate of State tax and refund would depend upon the State, which is not mentioned here.

Married Filing Jointly
Taxable Income Tax Rate
$0 – $18,650 10%
$18,651 – $75,900 $1,865 plus 15% of the amount over $18,650
$75,901 – $153,100 $10,452.50 plus 25% of the amount over $75,900
$153,101 – $233,350 $29,752.50 plus 28% of the amount over $153,100
$233,351 – $416,700 $52,222.50 plus 33% of the amount over $233,350
$416,701 – $470,700 $112,728 plus 35% of the amount over $416,700
$470,701 or more $131,628 plus 39.6% of the amount over $470,700

As per question wages is $65,000, it would fall in 2nd slab in bold. Tax is calculated as below:

=1865+ (65000-18651)*15%= 1865+ 6952.35= 8817.35

Tax Withheld = 15000, so refund would be 15,000- 8817.35= $6182.65